e^-3x=5
is the answer 5? Teachers examples weren't helpful!

Answer by edjones(8007)

e^-3x=5
ln(e^-3x)=ln(5)
-3x=ln(5)
x=ln(5)/-3
x=1.609/-3
x=-.536
.
Ed